OpenStudy (anonymous):

A waitress fills bowls with soup. She puts twothirds_large of a cup of soup in each bowl. How many bowls can the waitress prepare with 40 cups of soup?

OpenStudy (anonymous):

40 cups divided by 2/3 cups

OpenStudy (anonymous):

same concept that i keep repeating

OpenStudy (anonymous):

plz help

OpenStudy (dan815):

hello!

OpenStudy (anonymous):

hi

OpenStudy (dan815):

do u like to eat cookies

OpenStudy (anonymous):

yes

OpenStudy (dan815):

say you have a bag of cookies

OpenStudy (dan815):

there are 40 cookies in this bag

OpenStudy (dan815):

and you got friends over, and seeing as you are such a good friend, u want to share your cookies with them

OpenStudy (anonymous):

ok

OpenStudy (dan815):

so all your friends have 2 hands, and u put 1 cookie each hand of theirs

OpenStudy (dan815):

by the time you finish giving them all 2 cookies, ur done... your bag of cookies are done.. how many friends do you have?

OpenStudy (anonymous):

but do i get any cookies or gust them?

OpenStudy (dan815):

just them =[

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

20 friends

OpenStudy (dan815):

awesome!!

OpenStudy (dan815):

okay now lets do your question..

OpenStudy (dan815):

which has absolutely nothing to do with the cookie question, i just had to see if you can keep up with reading

OpenStudy (dan815):

lol

OpenStudy (anonymous):

ok

OpenStudy (dan815):

so lets see whats going on , when ever you see a question always visualize it

OpenStudy (dan815):

now you have 40 cups of soup!

OpenStudy (dan815):

and you can only pour 2/3rd of each cup into a bowl

OpenStudy (dan815):

how many bowls can you fill with 40 cups of soup!

OpenStudy (dan815):

we must see how many 2/3 rds are there in 40 cups

OpenStudy (dan815):

because 2/3rd = 1 bowl

OpenStudy (dan815):

the number of 2/3rds in 40 cups = number of bowl you will fill

OpenStudy (dan815):

can you solve this equation

OpenStudy (dan815):

have you done algebra yet

OpenStudy (anonymous):

yes but it is hard for me

OpenStudy (dan815):

okay nvm that algebra is needlely complicating then

OpenStudy (dan815):

look at this simple question

OpenStudy (dan815):

40 cookies and u give 2 cookies out to each person how many persons can u give to then

OpenStudy (anonymous):

20

OpenStudy (dan815):

okay what did u do in your mind there?

OpenStudy (dan815):

you said 2 * what = 40 right

OpenStudy (dan815):

or maybe you said 40/2 = what

OpenStudy (anonymous):

yes

OpenStudy (dan815):

thinking like this is the same 2*what=40 is the same thing as saying 40/what = 2 or 40/2=what?

OpenStudy (dan815):

thats all x really is lol! x is just that what!

OpenStudy (dan815):

anyway moving on now say you have 40 cookies, but u wanna give out 2/3rd of a cookie to each person, how many persons can u give to

OpenStudy (dan815):

remember how you did for 2 cookies a person 2*what=40 what = 20 or 40/2=what? what = 20 now lets do this for 2/3 2/3 * what = 40?? or 40/(2/3) = what??

OpenStudy (anonymous):

22...

OpenStudy (dan815):